# Hadamard-type formulas for real eigenvalues of canonically symplectic operators

### v2

@@ 1. Introduction
Recently many works have used the Maslov index to study nonlinear PDEs. In this paper, we are motivated by two recent works [CCLM23, LS20first]. The first discusses Hadamard-type formulas for rates of change of eigenvalues of self-adjoint realizations of elliptic operators on bounded domains and abstract symmetric operators in Hilbert spaces [LS20first]; the second discusses similar formulas for the rates of change of real eigenvalues, with respect to perturbations of the domain, of a canonically symplectic operator [KKS04] arising in the stability analysis of standing waves in the NLS equation on a compact interval subject to Dirichlet boundary conditions [CCLM23]. Here we extend the results of the latter by using the framework of the former: we study the variation of real eigenvalues of canonically symplectic operators where the off-diagonal blocks are self-adjoint extensions of densely-defined symmetric operators. Our results bridge the celebrated classical Rayleigh-Hadamard-Rellich theory for eigenvalue variation [Rayleigh45, hadamard1908, Rellich69] with the Arnold-Maslov-Keller index theory from symplectic geometry [Maslov65, A67, Keller58] for operators with the canonical symplectic structure. Since such operators are not self-adjoint, most of the monotonicity properties present in the self-adjoint setting are lost, leading to far richer spectral behaviour. Furthermore, as shown in [CCLM23] these local formulas can provide crucial information in determining global counts of eigenvalues. Before discussing our abstract setting and main results of the paper, we provide some examples, accompanied with numerical experiments, to motivate our analysis.
@@ 1.1. Motivating examples
The operator $\mathcal{N}$ is said to have the canonical symplectic structure [KKS04]. Its spectrum, consisting solely of eigenvalues due to the compactness of the domain, thus has four-fold symmetry in the complex plane. To determine the spectral stability of the standing wave, one must therefore determine the existence of any eigenvalues lying off the imaginary axis. While classical results [GSS87, Grill88, GSS90] are relevant in the present context, see also [Pel05, pelinovsky, KKS04, KKS05, KP12art], here we opt for a symplectic approach, which is related to the study of Hadamard-type variational formulas for parameter-dependent eigenvalues in an effort to further recent studies of the latter [CCLM23, LS20first]. Namely, we consider the problem on a family of shrinking subdomains, parameterised by $t\in(0,1]$, and study the spectrum of the problem on the full domain when $t=1$ by analysing the $t$-dependent spectrum when $\lambda=0$ and $t\in(0,1]$. This approach has been used to study the spectra of ordinary and partial differential operators in many works, for example, [Smale65, DJ11, JLM13, CJLS14, LSHad17].
@@
and because $\phi$ satisfies (1.5) we have $\operatorname{tr}\phi=(\Gamma_{0}\phi,\Gamma_{1}\phi)^{\top}\in\mathcal{L}$, hence $\phi\in\operatorname{dom}(\mathcal{A})$. We make the assumption that zero is a simple eigenvalue of $\mathcal{A}+F$; in the case of Neumann-Kirchhoff conditions (where $\alpha=0$), this assumption is generic with respect to the set of edge lengths $\{\ell_{i}\}$ [berk, Theorem 3.1.7]. We also assume invertibility of $\mathcal{A}+G$.
@@
It follows from the previous assumption that zero is a simple eigenvalue of $\mathcal{N}+V$ with eigenfunction $\boldsymbol{\phi}\coloneqq(0,\phi)^{\top}$. The real spectral index can then be deduced from an analysis of the eigenvalue curves. The following theorem extends the main result of [CCLM23] from the setting of a compact interval with Dirichlet boundary conditions to star graphs with $\delta$-type conditions on the central vertex and Dirichlet conditions at the free vertices.
@@
A corollary of Theorem 1.2 is the following Vakhitov-Kolokolov-type (VK) criterion [VK73, pelinovsky], which furnishes a convenient numerical tool – indeed one need only local data at $(\lambda,t)=(0,1)$ – to establish the existence of a positive real eigenvalue. Suppose $p_{c}=1$ and $q_{c}=0$. Let $t(\lambda)$ be the eigenvalue curve, given in Theorem 1.2, which passes through $(\lambda,t)=(0,1)$ and is defined for $0<|\lambda|\ll 1$. Suppose that $t^{\prime\prime}(\lambda)|_{\lambda=0}\neq 0$. Then $\mathcal{N}+V$ has a positive real eigenvalue if $t^{\prime\prime}(0)>0$, while $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ if $t^{\prime\prime}(0)<0$. The first part of this assertion, that is, existence of unstable eigenvalues follows from Theorem 1.2. While the second part can be shown as in [CCLM23, pelinovsky].
@@
where $a\approx 0.8660$ and $b=5,3,1$ in Fig. 1 (c), (d) and (e) respectively. These initial conditions are integrated forward on each edge under (1.4) until the first instance at which the functions are zero. Thus the waves satisfy (1.5) with $\alpha=0$ (i.e. Neumann-Kirchhoff conditions), are non-negative on $\mathcal{G}$ and all have one bump (non-monotonic profile) and two tails (monotonic profile) [PavaGoloshch18, Kairzhan19, DKP22]. Integrating along the orbits in the phase plane associated with (1.4) shows that in all cases we have $\ell_{2}\approx 0.262628$, while
@@
Using a homotopy argument and [berk, Theorem 5.2.8], one can show that $q_{c}=n_{-}(\mathcal{A}+F)=0$. Noting that the number of intersections of the eigenvalue curves with the line $\lambda=0$ for $t<1$ is the quantity $p_{c}+q_{c}$, Fig. 1 (c), (d) and (e) therefore numerically verify that the three different standing waves described above all satisfy the condition of Remark 1.5. If the eigenvalue curves are continuous and there are no points of horizontal tangency for $t<1$ away from $\lambda=0$, it follows that the existence of a positive real eigenvalue on the full domain (given by intersections with $t=1$) may be predicted from the concavity of the eigenvalue curve through the point $(\lambda,t)=(0,1)$: if $t^{\prime\prime}(\lambda)|_{\lambda=0}>0$ then $\mathcal{N}+V$ has a positive real eigenvalue, while if $t^{\prime\prime}(\lambda)|_{\lambda=0}<0$ then $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ [^2] It can be proven that there are no complex eigenvalues under the assumptions of Remark 1.5, see e.g. [CCLM23, pelinovsky]. Figure 1 therefore reflects a change in stability of the underlying wave as $b$ decreases (and the edge lengths $\ell_{1}$ and $\ell_{3}$ increase): the concavity of the eigenvalue curve changes from negative to positive, indicating that a pair of imaginary eigenvalues have bifurcated onto the real axis.
Following the analysis in [LS20first], in this paper we also concern ourselves with variations in the boundary conditions and their effect on the spectrum of canonically symplectic operators. To this end, in Fig. 2 we numerically compute the eigenvalue curves for three different operators $\mathcal{N}_{t}+V_{t}$ given by (1.11), but where the domain is a compact interval, i.e. $\operatorname{dom}A=H^{2}_{0}(0,\ell)$ and $\operatorname{dom}\mathcal{A}_{t}=\left\{u\in H^{2}:\operatorname{tr}_{t}u\in\mathcal{L}_{t}\right\}$, and:
@@ 1.2. Discussion of abstract setting and main results
Motivated by the above examples and following the analyses in [LS20first, CCLM23], in this paper we derive Hadamard formulas for the real eigenvalue curves of canonically symplectic operators, i.e. off-diagonal block operators where the blocks are self-adjoint extensions of densely-defined symmetric operators with finite deficiency indices. Here, as in [LS20first], we allow for $t$-dependence in the bounded perturbation of the operator, in the trace operator, and in the Lagrangian planes describing the boundary conditions. We adopt two different approaches: on the one hand, following [LS20first] we derive a symplectic resolvent difference formula which hinges on a modified Green’s identity for the operators of interest. We use these formulas to obtain a first order asymptotic expansion of the resolvent, which, in turn, we use to derive an Hadamard formula for the eigenvalue curve $\lambda(t)$. On the other hand, following [CCLM23] we use Lyaponuv-Schmidt reduction and the implicit function theorem to arrive at the same Hadamard-type formula for $\lambda(t)$, and to derive an additional formula for $t(\lambda)$. As outlined in the first example in the previous section, such a formula is necessary in order to determine global counts of eigenvalues. On a conceptual level, the main difference from the analysis in [CCLM23] is that we now allow for $t$-dependence in the Lagrangian plane describing the boundary conditions (where such Lagrangian plane was the (fixed) Dirichlet plane in [CCLM23]).
@@
Our third main result establishes the previous formula via a different method using Lyaponuv-Schmidt reduction and gives an expression for $t^{\prime}(\lambda)$ . In particular, if $\lambda_{0}\in\mathbb{R}$ is an isolated eigenvalue of $\mathcal{N}_{t_{0}}+V_{t_{0}}$ of geometric multiplicity $g$, we show in Proposition 4.6 that the set of points $(\lambda,t)$ near $(\lambda_{0},t_{0})$ such that $\ker(\mathcal{N}_{t}+V_{t}-\lambda)\neq\{0\}$ is given by the zero set of a $g\times g$ matrix $M(\lambda,t)$. We then compute Maslov crossing forms in Lemma 4.5 for the path of Lagrangian planes $\Upsilon_{\lambda,t}\coloneqq\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}$ on the intersection $\left(\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}\right)\cap\mathfrak{D}$, where $\mathbb{K}_{\lambda,t}\coloneqq\operatorname{T}_{t}(\ker((N^{*}+V_{t}^{*}-\lambda)\tau))$, $\mathcal{F}_{t}\coloneqq\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}$ and $\mathfrak{D}$ is the diagonal subspace in $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ (cf. (4.4)),
@@
The paper is organised as follows. In Section 2, we describe our abstract set-up and main assumptions, give the Green identity which underpins our analysis, and prove the new symplectic resolvent difference formula for canonically symplectic operators. In Section 3 we use our resolvent difference formula to give a first order expansion of the resolvent, and use this to prove an Hadamard formula for eigenvalue curves $\lambda(t)$. In Section 4 we compute expressions for the Maslov crossing forms, and use Lyaponuv–Schmidt reduction and the implicit function theorem to compute Hadamard formulas for the eigenvalue curves (namely for both $\lambda(t)$ and $t(\lambda)$). In Section 5 we use our expressions for the crossing forms to prove Theorem 1.2 for the example outlined in the introduction.
@@ 2. Set-up, symplectic resolvent difference formula and modified Green’s identity
Thus, under Hypothesis 2.1, the triple $(\mathfrak{H},\Gamma_{0},\Gamma_{1})$ is a boundary triplet for the adjoint operator $A^{*}$. Consequently, by [Behrndt_Hassi_deSnoo, Proposition 14.5], we have
@@ 3.1. Asymptotic expansions and a Hadamard formula
We note that (3.32) holds due to our assumption that $\lambda$ is a simple eigenvalue. It would still continue to hold for semi-simple eigenvalues (i.e., those with equal algebraic and geometric multiplicities). However in a more general case of non-semi-simple eigenvalues the right-hand side of (3.32) contains a nilpotent part, cf., e.g., [Kato, eq. (5.23) in Chapter I.5] which would enter an Hadamart-type formula for the derivative of eigenvalue curves.
@@
We remark that, since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, $P(t)$ and $S$ are not self-adjoint, see for example [Gohberg_Goldberg_Kaashoek, Proposition I.2.5]. However, using (2.17) and (3.5) we have
@@
Towards a Hadamard formula, we need to write down an asymptotic expansion for $P(t)H_{t}P(t)$ for $t$ near $t_{0}$. Since the domain of $P(t)H_{t}P(t)$ is $t$-dependent, we introduce the following transformation operators which allow us to work with a similar operator for which the domain is fixed, cf. [Kato, Section I.4.6]. To that end, we define $D(t)\coloneqq P(t)-P(t_{0})$ satisfying $\|D(t)\|_{\mathcal{B}(\mathcal{H})}=o(1)$ as $t\to t_{0}$, which follows from (3.21) and the definition of $P(t)$ (3.30). The following operators are then well defined for $t$ near $t_{0}$.
@@
Moreover, as in [Kato, Section I.4.6] and [F04, Proposition 2.18], we have that
@@
and that $U(t)$ maps $\operatorname{ran}P(t_{0})$ onto $\operatorname{ran}P(t)$ for $t$ near $t_{0}$. Unlike the analysis in [LS20first], since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, the Riesz projection $P(t)$ is not self-adjoint and therefore not orthogonal, and hence $U(t)$ is not unitary. (In fact, one can show that $U(t)^{-1}=\tau U(t)^{*}\tau$.) Nonetheless, only minor tweaks to the proof of [LS20first, Lemma 3.24] are required for the following lemma.
@@
Recalling that $U(t)$ is an isomorphism between $\operatorname{ran}P(t_{0})$ and $\operatorname{ran}P(t)$, see [Kato, Section I.4.6], [F04, Proposition 2.18], we note that $(\mathcal{N}_{t}+V_{t})|_{\operatorname{ran}P(t)}$ is similar to
@@
for $t$ near $t_{0}$. Thus the eigenvalues of these operators coincide for all $t$ near $t_{0}$, see [Kato, Section I.5.7]. Asymptotically expanding the eigenvalues of the latter using the finite dimensional first order perturbation theory, specifically [Kato, Theorem II.5.11], we deduce (3.53).
@@
4. Hadamard formulas via Lyaponuv-Schmidt
In this section, under certain assumptions we prove the existence of eigenvalue curves $\lambda(t)\in\operatorname{Spec}\left(\mathcal{N}_{t}+V_{t}\right)$ or $\lambda\in\operatorname{Spec}\left(\mathcal{N}_{t(\lambda)}+V_{t(\lambda)}\right)$, and compute local formulas for their derivatives. The analysis is similar to that in [CCLM23], however the situation herein differs due to the $t$-dependence of the Lagrangian planes describing the boundary conditions.

### v3

@@ 1. Introduction
Recently many works have used the Maslov index to study nonlinear PDEs. In this paper, we are motivated by two recent works [CCL+23, LS24]. The first discusses Hadamard-type formulas for rates of change of eigenvalues of self-adjoint realizations of elliptic operators on bounded domains and abstract symmetric operators in Hilbert spaces [LS24]; the second discusses similar formulas for the rates of change of real eigenvalues, with respect to perturbations of the domain, of a canonically symplectic operator [KKS04] arising in the stability analysis of standing waves in the NLS equation on a compact interval subject to Dirichlet boundary conditions [CCL+23]. Here we extend the results of the latter by using the framework of the former: we study the variation of real eigenvalues of canonically symplectic operators where the off-diagonal blocks are self-adjoint extensions of densely-defined symmetric operators. Our results bridge the celebrated classical Rayleigh-Hadamard-Rellich theory for eigenvalue variation [RAY45, HAD68, REL69] with the Arnold-Maslov-Keller index theory from symplectic geometry [MAS65, ARN67, KEL58] for operators with the canonical symplectic structure. Since such operators are not self-adjoint, most of the monotonicity properties present in the self-adjoint setting are lost, leading to far richer spectral behaviour. Furthermore, as shown in [CCL+23] these local formulas can provide crucial information in determining global counts of eigenvalues. Before discussing our abstract setting and main results of the paper, we provide some examples, accompanied with numerical experiments, to motivate our analysis.
@@ 1.1. Motivating examples
The operator $\mathcal{N}$ is said to have the canonical symplectic structure [KKS04]. Its spectrum, consisting solely of eigenvalues due to the compactness of the domain, thus has four-fold symmetry in the complex plane. To determine the spectral stability of the standing wave, one must therefore determine the existence of any eigenvalues lying off the imaginary axis. While classical results [GSS87, GRI88, GSS90] are relevant in the present context, see also [PEL05, PEL11, KKS04, KKS05, KP12], here we opt for a symplectic approach, which is related to the study of Hadamard-type variational formulas for parameter-dependent eigenvalues in an effort to further recent studies of the latter [CCL+23, LS24]. Namely, we consider the problem on a family of shrinking subdomains, parameterised by $t\in(0,1]$, and study the spectrum of the problem on the full domain when $t=1$ by analysing the $t$-dependent spectrum when $\lambda=0$ and $t\in(0,1]$. This approach has been used to study the spectra of ordinary and partial differential operators in many works, for example, [SMA65, DJ11, JLM13, CJL+16, LS17].
@@
and because $\phi$ satisfies (1.5) we have $\operatorname{tr}\phi=(\Gamma_{0}\phi,\Gamma_{1}\phi)^{\top}\in\mathcal{L}$, hence $\phi\in\operatorname{dom}(\mathcal{A})$. We make the assumption that zero is a simple eigenvalue of $\mathcal{A}+F$; in the case of Neumann-Kirchhoff conditions (where $\alpha=0$), this assumption is generic with respect to the set of edge lengths $\{\ell_{i}\}$ [BK13, Theorem 3.1.7]. We also assume invertibility of $\mathcal{A}+G$.
@@
It follows from the previous assumption that zero is a simple eigenvalue of $\mathcal{N}+V$ with eigenfunction $\boldsymbol{\phi}\coloneqq(0,\phi)^{\top}$. The real spectral index can then be deduced from an analysis of the eigenvalue curves. The following theorem extends the main result of [CCL+23] from the setting of a compact interval with Dirichlet boundary conditions to star graphs with $\delta$-type conditions on the central vertex and Dirichlet conditions at the free vertices.
@@
A corollary of Theorem 1.2 is the following Vakhitov-Kolokolov-type (VK) criterion [VK73, PEL11], which furnishes a convenient numerical tool – indeed one need only local data at $(\lambda,t)=(0,1)$ – to establish the existence of a positive real eigenvalue. Suppose $p_{c}=1$ and $q_{c}=0$. Let $t(\lambda)$ be the eigenvalue curve, given in Theorem 1.2, which passes through $(\lambda,t)=(0,1)$ and is defined for $0<|\lambda|\ll 1$. Suppose that $t^{\prime\prime}(\lambda)|_{\lambda=0}\neq 0$. Then $\mathcal{N}+V$ has a positive real eigenvalue if $t^{\prime\prime}(0)>0$, while $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ if $t^{\prime\prime}(0)<0$. The first part of this assertion, that is, existence of unstable eigenvalues follows from Theorem 1.2. While the second part can be shown as in [CCL+23, PEL11].
@@
where $a\approx 0.8660$ and $b=5,3,1$ in Fig. 1 (c), (d) and (e) respectively. These initial conditions are integrated forward on each edge under (1.4) until the first instance at which the functions are zero. Thus the waves satisfy (1.5) with $\alpha=0$ (i.e. Neumann-Kirchhoff conditions), are non-negative on $\mathcal{G}$ and all have one bump (non-monotonic profile) and two tails (monotonic profile) [AG18, KAI19, KNP22]. Integrating along the orbits in the phase plane associated with (1.4) shows that in all cases we have $\ell_{2}\approx 0.262628$, while
@@
Using a homotopy argument and [BK13, Theorem 5.2.8], one can show that $q_{c}=n_{-}(\mathcal{A}+F)=0$. Noting that the number of intersections of the eigenvalue curves with the line $\lambda=0$ for $t<1$ is the quantity $p_{c}+q_{c}$, Fig. 1 (c), (d) and (e) therefore numerically verify that the three different standing waves described above all satisfy the condition of Remark 1.5. If the eigenvalue curves are continuous and there are no points of horizontal tangency for $t<1$ away from $\lambda=0$, it follows that the existence of a positive real eigenvalue on the full domain (given by intersections with $t=1$) may be predicted from the concavity of the eigenvalue curve through the point $(\lambda,t)=(0,1)$: if $t^{\prime\prime}(\lambda)|_{\lambda=0}>0$ then $\mathcal{N}+V$ has a positive real eigenvalue, while if $t^{\prime\prime}(\lambda)|_{\lambda=0}<0$ then $\operatorname{Spec}(\mathcal{N}+V)\subset i\mathbb{R}$ [^2] It can be proven that there are no complex eigenvalues under the assumptions of Remark 1.5, see e.g. [CCL+23, PEL11]. Figure 1 therefore reflects a change in stability of the underlying wave as $b$ decreases (and the edge lengths $\ell_{1}$ and $\ell_{3}$ increase): the concavity of the eigenvalue curve changes from negative to positive, indicating that a pair of imaginary eigenvalues have bifurcated onto the real axis.
Following the analysis in [LS24], in this paper we also concern ourselves with variations in the boundary conditions and their effect on the spectrum of canonically symplectic operators. To this end, in Fig. 2 we numerically compute the eigenvalue curves for three different operators $\mathcal{N}_{t}+V_{t}$ given by (1.11), but where the domain is a compact interval, i.e. $\operatorname{dom}A=H^{2}_{0}(0,\ell)$ and $\operatorname{dom}\mathcal{A}_{t}=\left\{u\in H^{2}:\operatorname{tr}_{t}u\in\mathcal{L}_{t}\right\}$, and:
@@ 1.2. Discussion of abstract setting and main results
Motivated by the above examples and following the analyses in [LS24, CCL+23], in this paper we derive Hadamard formulas for the real eigenvalue curves of canonically symplectic operators, i.e. off-diagonal block operators where the blocks are self-adjoint extensions of densely-defined symmetric operators with finite deficiency indices. Here, as in [LS24], we allow for $t$-dependence in the bounded perturbation of the operator, in the trace operator, and in the Lagrangian planes describing the boundary conditions. We adopt two different approaches: on the one hand, following [LS24] we derive a symplectic resolvent difference formula which hinges on a modified Green’s identity for the operators of interest. We use these formulas to obtain a first order asymptotic expansion of the resolvent, which, in turn, we use to derive an Hadamard formula for the eigenvalue curve $\lambda(t)$. On the other hand, following [CCL+23] we use Lyapunov-Schmidt reduction and the implicit function theorem to arrive at the same Hadamard-type formula for $\lambda(t)$, and to derive an additional formula for $t(\lambda)$. As outlined in the first example in the previous section, such a formula is necessary in order to determine global counts of eigenvalues. On a conceptual level, the main difference from the analysis in [CCL+23] is that we now allow for $t$-dependence in the Lagrangian plane describing the boundary conditions (where such Lagrangian plane was the (fixed) Dirichlet plane in [CCL+23]).
@@
Our third main result establishes the previous formula via a different method using Lyapunov-Schmidt reduction and gives an expression for $t^{\prime}(\lambda)$ . In particular, if $\lambda_{0}\in\mathbb{R}$ is an isolated eigenvalue of $\mathcal{N}_{t_{0}}+V_{t_{0}}$ of geometric multiplicity $g$, we show in Proposition 4.6 that the set of points $(\lambda,t)$ near $(\lambda_{0},t_{0})$ such that $\ker(\mathcal{N}_{t}+V_{t}-\lambda)\neq\{0\}$ is given by the zero set of a $g\times g$ matrix $M(\lambda,t)$. We then compute Maslov crossing forms in Lemma 4.5 for the path of Lagrangian planes $\Upsilon_{\lambda,t}\coloneqq\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}$ on the intersection $\left(\mathbb{K}_{\lambda,t}\oplus\mathcal{F}_{t}\right)\cap\mathfrak{D}$, where $\mathbb{K}_{\lambda,t}\coloneqq\operatorname{T}_{t}(\ker((N^{*}+V_{t}^{*}-\lambda)\tau))$, $\mathcal{F}_{t}\coloneqq\operatorname{ran}P_{t}\times\operatorname{ran}Q_{t}$ and $\mathfrak{D}$ is the diagonal subspace in $\mathfrak{H}^{4}\times\mathfrak{H}^{4}$ (cf. (4.4)),
@@
The paper is organised as follows. In Section 2, we describe our abstract set-up and main assumptions, give the Green identity which underpins our analysis, and prove the new symplectic resolvent difference formula for canonically symplectic operators. In Section 3 we use our resolvent difference formula to give a first order expansion of the resolvent, and use this to prove an Hadamard formula for eigenvalue curves $\lambda(t)$. In Section 4 we compute expressions for the Maslov crossing forms, and use Lyapunov–Schmidt reduction and the implicit function theorem to compute Hadamard formulas for the eigenvalue curves (namely for both $\lambda(t)$ and $t(\lambda)$). In Section 5 we use our expressions for the crossing forms to prove Theorem 1.2 for the example outlined in the introduction.
@@ 2. Set-up, symplectic resolvent difference formula and modified Green’s identity
Thus, under Hypothesis 2.1, the triple $(\mathfrak{H},\Gamma_{0},\Gamma_{1})$ is a boundary triplet for the adjoint operator $A^{*}$. Consequently, by [BHd02, Proposition 14.5], we have
@@ 3.1. Asymptotic expansions and a Hadamard formula
We note that (3.32) holds due to our assumption that $\lambda$ is a simple eigenvalue. It would still continue to hold for semi-simple eigenvalues (i.e., those with equal algebraic and geometric multiplicities). However in a more general case of non-semi-simple eigenvalues the right-hand side of (3.32) contains a nilpotent part, cf., e.g., [KAT80, eq. (5.23) in Chapter I.5] which would enter an Hadamard-type formula for the derivative of eigenvalue curves.
@@
We remark that, since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, $P(t)$ and $S$ are not self-adjoint, see for example [GGK90, Proposition I.2.5]. However, using (2.17) and (3.5) we have
@@
Towards a Hadamard formula, we need to write down an asymptotic expansion for $P(t)H_{t}P(t)$ for $t$ near $t_{0}$. Since the domain of $P(t)H_{t}P(t)$ is $t$-dependent, we introduce the following transformation operators which allow us to work with a similar operator for which the domain is fixed, cf. [KAT80, Section I.4.6]. To that end, we define $D(t)\coloneqq P(t)-P(t_{0})$ satisfying $\|D(t)\|_{\mathcal{B}(\mathcal{H})}=o(1)$ as $t\to t_{0}$, which follows from (3.21) and the definition of $P(t)$ (3.30). The following operators are then well defined for $t$ near $t_{0}$.
@@
Moreover, as in [KAT80, Section I.4.6] and [FUR04, Proposition 2.18], we have that
@@
and that $U(t)$ maps $\operatorname{ran}P(t_{0})$ onto $\operatorname{ran}P(t)$ for $t$ near $t_{0}$. Unlike the analysis in [LS24], since $\mathcal{N}_{t}+V_{t}$ is not self-adjoint, the Riesz projection $P(t)$ is not self-adjoint and therefore not orthogonal, and hence $U(t)$ is not unitary. (In fact, one can show that $U(t)^{-1}=\tau U(t)^{*}\tau$.) Nonetheless, only minor tweaks to the proof of [LS24, Lemma 3.24] are required for the following lemma.
@@
Recalling that $U(t)$ is an isomorphism between $\operatorname{ran}P(t_{0})$ and $\operatorname{ran}P(t)$, see [KAT80, Section I.4.6], [FUR04, Proposition 2.18], we note that $(\mathcal{N}_{t}+V_{t})|_{\operatorname{ran}P(t)}$ is similar to
@@
for $t$ near $t_{0}$. Thus the eigenvalues of these operators coincide for all $t$ near $t_{0}$, see [KAT80, Section I.5.7]. Asymptotically expanding the eigenvalues of the latter using the finite dimensional first order perturbation theory, specifically [KAT80, Theorem II.5.11], we deduce (3.53).
@@
4. Hadamard formulas via Lyapunov-Schmidt
In this section, under certain assumptions we prove the existence of eigenvalue curves $\lambda(t)\in\operatorname{Spec}\left(\mathcal{N}_{t}+V_{t}\right)$ or $\lambda\in\operatorname{Spec}\left(\mathcal{N}_{t(\lambda)}+V_{t(\lambda)}\right)$, and compute local formulas for their derivatives. The analysis is similar to that in [CCL+23], however the situation herein differs due to the $t$-dependence of the Lagrangian planes describing the boundary conditions.